OpenStudy (anonymous):

Question: Take the limit to infinity of (3x-5x^2)/(2x^2 +4)

OpenStudy (anonymous):

Is it -infinity/infinity?

OpenStudy (whpalmer4):

that's the limit as x -> infinity?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lim x->inf

OpenStudy (whpalmer4):

no, then with the highest exponent of x in both numerator and denominator being equal, it becomes just the ratio of their coefficients

OpenStudy (whpalmer4):

you agree that the behavior of the term with the largest exponent dominates the behavior as x->infinity, right?

OpenStudy (anonymous):

yes, I remember that, just not sure how to carry it out

OpenStudy (whpalmer4):

okay, so "simplify" to \[\frac{-5x^2}{2x^2}\]what's that going to do as \(x\rightarrow\infty\)?

OpenStudy (anonymous):

a negative number

OpenStudy (whpalmer4):

it's going to have the same value it has at any non-zero value of \(x\), which is...

OpenStudy (anonymous):

-5/2 ?

OpenStudy (whpalmer4):

indeed

OpenStudy (anonymous):

ah ok, so when I'm faced with these problems

OpenStudy (anonymous):

drop everything but the highest power of x in the numerator and denominator

OpenStudy (anonymous):

then determine it's simplest form?

OpenStudy (whpalmer4):

if the numerator has a term that has a higher exponent, then you just find the limit of that term. \[\lim_{ x\rightarrow\infty} \frac{x^2}{3x} = \infty\]because there's nothing "holding it back" \[\lim_{x\rightarrow\infty}\frac{x}{3x^2} = 0\] it's really only the case where the highest exponent in both numerator and denominator are equal that anything interesting happens...

OpenStudy (anonymous):

ah ok

OpenStudy (whpalmer4):

Yeah, it's just like finding the end behavior of a polynomial. Only the highest exponent term has any real impact once you move away from 0.

OpenStudy (whpalmer4):

One of the relatively few times you get to just throw stuff away and still get the right answer :-)

OpenStudy (anonymous):

XD thanks again

OpenStudy (whpalmer4):

you bet